OpenStudy (anonymous):

f(x)=(7x+12)^6 find the second derivative f"(x).

OpenStudy (akshay_budhkar):

f'(x)=6(7x+12)^5(7)\ f''(x)=6(5)(7x+12)^4(7)(7)

OpenStudy (anonymous):

so who's right?

OpenStudy (akshay_budhkar):

both are same :P

OpenStudy (anonymous):

f'(x) = 42 (7x + 12) ^5 f"(x) = 1470 (7x + 12) ^4

OpenStudy (akshay_budhkar):

ok i was right :P

OpenStudy (anonymous):

ok cause i was getting 1470

OpenStudy (anonymous):

Yep! I'm trying to multiply it out!

OpenStudy (anonymous):

when i multiplied 30 times 7 times 7

OpenStudy (anonymous):

thanks for helping :)

OpenStudy (anonymous):

We're all trying to help :)